1.Given below are the number of successes and sample size for a simple random sample from a population. x​= 7, n = 50​, ​98% level. the 98% confidence interval is from _ to _.
Answer #1

The 98% confidence interval for p:

The formula for the CI on parameter p is:
CI on parameter p


The unbiased point estimator, p is the proportion of “successes” in a Bernoulli trial. As a formula, that’s:
unbiased point estimator

\widehat{p}=\frac{7}{50}=0.14

Zα/2 = 2.326 (from z table)

Therefore, the 98% confidence interval for p:

0.14-2.326\times \sqrt{\frac{0.14(1-0.14)}{50}}<p<0.14+2.326\times \sqrt{\frac{0.14(1-0.14)}{50}}

0.14-0.049<p<0.14+0.049

0.091<p<0.189

The 98% confidence interval is from 0.091 to 0.189